OpenStudy (annie96):

A collection of quarters and nickels is worth $3.70. There are 22 coins in all. How many of each coin are there. A) 13nickels 9dimes B)9nickels 13dimes C) 11nickels 11dimes D)10nickels 12dimes

OpenStudy (anonymous):

Q+N=22 Q=22-N 25Q+5N=370 NOW SUBSTITUTE (22-N) FOR N IN THE SECOND EQUATION THUS: 25(22-N)+5N=370 550-25N+5N=370 -20N=370-550 -20N=-180 N=-180/-20 N=9 NUMBER OF NICKELS Q+9=22 Q=22-9 Q=13 NUMBER OF QUARTERS PROOF 25*13+5*9=370 325+45=370 370=370

OpenStudy (annie96):

So the answer is B.